likeabossssssss (likeabossssssss):

Which shows the dimensions of two rectangular prisms that have volumes of 240 ft3 but different surface areas? A. 8 ft by 5 ft by 6 ft; 5 ft by 6 ft by 8 ft B. 8 ft by 5 ft by 6 ft; 10 ft by 8 ft by 3 ft C. 15 ft by 8 ft by 2 ft; 6 ft by 4 ft by 15 ft D. 12 ft by 6 ft by 5 ft; 8 ft by 10 ft by 3 ft

likeabossssssss (likeabossssssss):

I think it is d

OpenStudy (dayakar):

check using v= l*b*h and surface area = 2(lb +bh+lh)

likeabossssssss (likeabossssssss):

ok thx

OpenStudy (dayakar):

:)